# Highly efficient quantum Stirling engine using multilayer Graphene

### v2

@@ I Introduction
Quantum thermodynamics examines how the fundamental principles of energy transfer, heat exchange, and entropy apply to small systems where quantum effects such as coherence, entanglement, spin correlations, and discrete energy levels give rise to thermodynamical properties with no classical counterpart, and where quantum statistics provides a natural framework to describe these phenomena Vinjanampathy and Anders (2016); Kosloff (2013); Deffner and Campbell (2019); Millen and Xuereb (2016); Binder et al. (2015, 2018); Myers et al. (2022); Zanin et al. (2019); Vieira et al. (2023); De Chiara et al. (2018); Castorene et al. (2025a); Liuzzo-Scorpo et al. (2016); Brandao et al. (2015).
@@
In this context, considerable attention has been devoted to understanding how quantum systems, such as trapped ions, superconducting qubits, quantum dots Esposito et al. (2010); Erdman et al. (2017); Josefsson et al. (2018, 2019); Josefsson and Leijnse (2020); Du et al. (2020), and low-dimensional electronic materials and their diverse behaviors can be harnessed to implement heat engines capable of producing useful work. Among quantum thermal machines, the Stirling cycle has attracted particular interest due to its potential to achieve high efficiency in regimes where thermalization and isochoric processes can be controlled with precision Saygin and Sisman (2001); Sisman and Saygin (2001); Cruz et al. (2023); Raja et al. (2021); Cakmak and Sedehi (2023); Castorene et al. (2024); Yin et al. (2018); Papadatos (2023); Wu et al. (1998); Rastegar-Sedehi et al. (2025); Huang et al. (2002); Xia et al. (2024); Li et al. (2025); Araya et al. (2023); Chatterjee et al. (2021); Zhang et al. (2024); Koyanagi and Tanimura (2024). In light of these considerations, theoretical and experimental investigations have shown that quantum heat engines based on discrete energy spectra may surpass their classical counterparts in specific regimes, particularly when the working substance exhibits a tunable energy level structure and distinctive statistical properties Peña et al. (2019); Deffner (2018).
@@
Graphene and its multilayer stacks, when subjected to perpendicular magnetic fields Castro Neto et al. (2009); McCann and Fal’ko (2006); Min and MacDonald (2008); Yuan et al. (2011); Goerbig (2011) stand out as promising platforms for quantum heat engines. Beyond their ease of fabrication and high degree of experimental control Koshino and Ando (2008); Nakamura and Hirasawa (2008); Ghai et al. (2024); Riffo et al. (2024); Rickhaus et al. (2012); Jahani et al. (2023); Valenzuela et al. (2015); Peng (2011); Bao et al. (2011), manifestations of their quantized spectrum are robust against temperature, since these systems exhibit quantum Hall plateaus even at room temperatures under sufficiently high magnetic fields Novoselov et al. (2007) – a critical feature for practical applications. Furthermore, the large degeneracy of the Landau levels involves a macroscopic number of participating electrons, $N$, even in micron-size flakes Goerbig (2011) This provides a unique framework to explore quantum corrections to thermodynamic quantities stemming from the discrete Landau-level spectrum, while remaining within the thermodynamic limit. Consequently, the small-$N$ thermal fluctuations typically found in few-body systems (both classical and quantum) are suppressed, making it possible to isolate genuine quantum effects. Very importantly, when driven by a varying magnetic field $B$, these quantum systems - unlike most others - avoid the dissipation associated to Landau-Zener transitions Kayanuma and Nakayama (1998); Dai et al. (2025), as Landau levels do not cross with changes in $B$. Finally, like other two-dimensional materials, the chemical potential (or doping) can be precisely tuned via electrostatic gating, offering a robust mechanism for controlling the engine’s working fluid.
@@
Several studies have proposed graphene-based systems, including quantum dots and multilayer structures, as promising candidates for quantum engines Mani et al. (2019); Singh and Benjamin (2021); Myers et al. (2023); Arroyo-Gascon et al. (2023); Kalugin et al. (2017); Pelc et al. (2015); Suarez Morell and Foa Torres (2012); Soufy and Benjamin (2025). While these pioneering works offer valuable insights, they predominantly rely on Boltzmann statistics Peña et al. (2020); Myers et al. (2023, 2022); Singh and Benjamin (2021). By neglecting the fermionic nature of the electronic working fluid and assuming charge neutrality, the lack of quantitative accuracy of their results and the narrow focus of their setting call for a reexamination.
@@
This study examines a quantum Stirling cycle implemented in monolayer, AB-stacked bilayer, and ABC-stacked trilayer graphene under perpendicular magnetic fields as the working medium. Employing a low-energy Landau-level description and Fermi–Dirac statistics at finite doping, we correct the previous studies, expand they scope beyond the charge-neutrality point, and provide a thermodynamic analysis more closely aligned with experimentally accessible conditions Pan et al. (2016); Moore et al. (2025). Within a grand-canonical framework, we evaluate key observables such as internal energy and entropy while explicitly accounting for the degeneracies specific to each multilayer configuration.
@@ II.1 Landau Level Spectrum in Multilayer Graphene
At low electronic doping, the low-energy LL spectra of mono-, AB-bilayer, and ABC-trilayer graphene in a perpendicular magnetic field (as shown in Fig. 1) can be cast into a single compact form that highlights the multilayer scalings McCann and Fal’ko (2006); Min and MacDonald (2008); Yuan et al. (2011); Castro Neto et al. (2009); Bao et al. (2011).
@@
where $n$ is the LL index and $\pm$ indicate conduction and valence band LLs. The $J$-dependent cyclotron frequency $\omega_{c}^{(J)}$ encodes the microscopic constants of each system. Equation (1) immediately yields the characteristic magnetic-field scaling
@@
For the three cases addressed here, the corresponding prefactors are:
@@
$\omega_{c}^{(1)}$ $=\sqrt{\frac{2ev_{F}^{2}B}{\hbar}},$
@@
$\omega_{c}^{(2)}$ $=\frac{eB}{m^{*}},\qquad m^{*}=\frac{t_{\perp}}{2v_{F}^{2}},$
@@
$\omega_{c}^{(3)}$ $=\frac{\left(2eBv_{F}^{2}\right)^{3/2}\sqrt{\hbar}}{t_{\perp}^{2}}.$
@@ III.1 Grand canonical formalism
In recent studies using graphene as a thermal machine, the thermodynamic behavior of graphene multilayers has been investigated through thermal states based on the canonical ensemble and Landau-level spectra Soufy and Benjamin (2025); Myers et al. (2023), neglecting the fermionic nature of the working fluid. In contrast, in this manuscript the thermodynamic properties of electrons in graphene multilayer systems under perpendicular magnetic fields are naturally described within the grand-canonical ensemble using Fermi–Dirac statistics, where both the temperature $T$ and the chemical potential $\mu$ serve as control parameters.
@@
Figure 2: Color map of the total number of electrons $\expectationvalue{N}$ as a function of magnetic field B[T]B\,[T] and chemical potential $\mu\$[meV] for (a) monolayer, (b) bilayer and (c) trilayered graphene system under a fixed temperature $T=50$ K.
For multilayer graphene with quantized Landau levels $\varepsilon_{n,\pm}^{(J)}$ (with $J=1,2,3$ denoting monolayer, bilayer, and trilayer), the grand-canonical partition function can be written explicitly as
@@
$S(T,B,\mu)$ $=-k_{B}\int\limits_{-\infty}^{\infty}d\varepsilon\,\nu^{(J)}(\varepsilon)\,\left[f\ln f+(1-f)\ln(1-f)\right].$ (13)
Both approaches are formally equivalent: the discrete-level formulation based on the grand potential and the continuum representation using the DOS combined with the Fermi–Dirac distribution, both yielding to identical thermodynamic quantities. It is worth emphasizing that the DOS representation makes explicit the role of Landau-level degeneracies, particularly the zero-energy states in all systems. We adopt the latter approach for our calculations. To guarantee numerical stability across the entire parameter space, the sums over Landau levels are truncated at a cutoff index of $n=2000$, or when the relative difference between consecutive evaluations falls below $10^{-6}$. Crucially, for the specific temperature and magnetic field regimes considered, thermal population is confined to the lowest levels, with contributions from $n>20$ being exponentially suppressed. Consequently, the effective number of participating states is limited to $n\lesssim 20$, which is fully consistent with the $\delta$-function DOS representation employed in this work Vidarte and Lewenkopf (2022).
@@ III.2 Energy Scales and Parametrization of the Chemical Potential
Having established the theoretical framework, we now estimate the order of magnitude of the energy separation between the ground state and the first accessible Landau excitation as the magnetic field $B$ is varied. This quantity ($\Delta$) sets the relevant energy scale for the thermodynamic analysis, since it determines the temperatures required to thermally populate excited states as the field is varied Castro Neto et al. (2009); McCann and Fal’ko (2006); Bao et al. (2011); Min and MacDonald (2008).
In monolayer graphene, the spacing between the ground and first excited Landau levels can be expressed as
$\Delta_{\mathrm{mono}}(B)=\hbar\omega_{c}^{(1)}\simeq 36.3\,\sqrt{B}\,\quantity[\frac{\mathrm{meV}}{\sqrt{\mathrm{T}}}].$ (14)
yielding $\Delta_{\mathrm{mono}}(1\,\mathrm{T})\approx 36.3\,\mathrm{meV}$, $\Delta_{\mathrm{mono}}(3.11\,\mathrm{T})\approx 64.0\,\mathrm{meV}$, and $\Delta_{\mathrm{mono}}(5\,\mathrm{T})\approx 81.2\,\mathrm{meV}$.
In AB–stacked bilayer graphene, described at low energies by a parabolic band with effective mass $m^{*}$, the first nonzero Landau level appears at $n=2$. The corresponding separation is
$\Delta_{\mathrm{bi}}(B)=\hbar\omega_{c}^{(2)}\simeq 4.77\times B\,\quantity[\frac{\mathrm{meV}}{\mathrm{T}}].$ (15)
which gives $\Delta_{\mathrm{bi}}(1\,\mathrm{T})\approx 4.77\,\mathrm{meV}$, $\Delta_{\mathrm{bi}}(3.11\,\mathrm{T})\approx 14.8\,\mathrm{meV}$, and $\Delta_{\mathrm{bi}}(5\,\mathrm{T})\approx 23.9\,\mathrm{meV}$.
For rhombohedral (ABC) trilayer graphene, the first accessible excitation occurs at $n=3$, with
[…]
leading to $\Delta_{\mathrm{tri}}(1\,\mathrm{T})\approx 0.77\,\mathrm{meV}$, $\Delta_{\mathrm{tri}}(3.11\,\mathrm{T})\approx 4.23\,\mathrm{meV}$, and $\Delta_{\mathrm{tri}}(5\,\mathrm{T})\approx 8.62\,\mathrm{meV}$.
For reference to experimental scales, we provide an estimate obtained from the magnetic length, $\ell_{B}=\sqrt{\hbar/(eB)}$ Castro Neto et al. (2009); Goerbig (2011). By defining $N_{e}$ as the number of electrons within the magnetic length area $\pi\ell_{B}^{2}$, the sheet density can be written as
@@
Fig. 2 provides an overview of the relation between the magnetic field $B$, the chemical potential $\mu$, and the resulting particle number $\langle N\rangle$ for mono-, bi-, and trilayer graphene at fixed temperature of $T=50$ K . In this figure, $N$ is not fixed; instead, we map the full surface $\langle N(50,B,\mu)\rangle$ in order to illustrate how the distinct Landau–level structures of each multilayer system shape the dependence of the particle number on $\mu$ and $B$. This exploration helps visualize the scaling behaviors originating from the linear, quadratic, and cubic dispersions of the three stackings.
@@ III.3 Thermodynamic Properties: Internal Energy and Entropy of Multilayer Graphene
The next discussion focuses on the entropy to further characterize the thermodynamic response. Figure 4 presents the temperature dependence of the entropy $S(T,B,\mu(T,B))$} for monolayer, bilayer, and trilayer graphene at different magnetic fields, with the particle number fixed at $N_{e}=15$. While all three systems exhibit a monotonic increase of entropy with temperature, notable differences arise in magnitude, curvature, and sensitivity to the magnetic field, underscoring the distinct thermodynamic responses of each multilayer configuration. In Fig. 4(a), the entropy of monolayer graphene remains negligible at low temperatures and begins to rise appreciably only above $T\simeq 25\,\mathrm{K}$, where thermal excitations populate higher Landau levels. The increase is markedly nonlinear, particularly at lower magnetic fields. This behavior reflects the relativistic Landau-level structure where stronger magnetic fields enlarge the spacing between energy levels and thereby suppress thermal activation. Consequently, higher fields substantially reduce the entropy and produce a clear separation among the curves.
@@ III.3.1 Quantum Stirling Cycle
Beyond these two classical regimes, previous studies have shown that the quantum Stirling cycle may also operate in two additional regimes that arise purely from its quantum character Rastegar-Sedehi and Cruz (2024); Castorene et al. (2025b); Rastegar-Sedehi and Cruz (2025); Castorene et al. (2024); Araya et al. (2023); Behrendt and Deffner (2025). In the accelerator regime, both the total work and the heat released to the cold reservoir are negative, indicating that external work input is accompanied by simultaneous heating of both reservoirs. In the heater regime, all heat flows are negative and $W<0$, so that the external work is entirely dissipated as heat into both reservoirs.
Importantly, these operational regimes are not exclusive to intrinsic properties of stacked graphene systems. It has been shown that twisted bilayer graphene, when employed as a Stirling engine at temperatures within 50 to 150 K, can also exhibit this diversity of thermodynamic behaviors Soufy and Benjamin (2025). As will be shown in later sections, the exotic thermodynamic properties of the system, which arise from the constraint of particle number conservation, can induce these additional operational regimes.
@@ IV.1 Stirling Engine Efficiency
For the bilayer graphene in Fig. 6(b), the system operates as an engine across the entire $B_{L}$ range considered, with the high magnetic field fixed at $B_{H}=3.18\,\mathrm{T}$. The thermal reservoirs are maintained at $T_{L}=45.9\,\mathrm{K}$ and $T_{H}=47.7\,\mathrm{K}$, corresponding to a Carnot efficiency of $\eta_{C}\simeq 0.04$. Under these conditions, the curves for $N_{e}=5$ and $N_{e}=15$ attain the Carnot bound at finite values of $B_{L}$, with the $N_{e}=15$ case attaining $\eta_{C}$ at lower $B_{L}$ values than $N_{e}=5$. In contrast, the $N_{e}=8$ curve remains slightly below $\eta_{C}$ over the full range, exhibiting a broad maximum just under the Carnot line at $B_{L}\simeq 0.8$[T]. This reflects a more balanced competition between the isothermal and isomagnetic contributions arising from the parametrization of the chemical potential. This intermediate behavior highlights how the bilayer spectrum, with its linear scaling in $B$, provides a broader operational window: depending on the carrier number, the cycle can either saturate the Carnot limit or remain close to it with finite work output. This behavior contrasts with results for twisted bilayer graphene studied within the canonical ensemble Soufy and Benjamin (2025). While such systems can reach Carnot efficiency in Otto cycles Singh and Benjamin (2021) and exhibit high efficiencies in quantum Stirling cycles, saturation of the Carnot bound in the Stirling cycle has not been reported in that framework. This suggests that a Fermi–Dirac description with fixed particle number enhances thermodynamic performance in graphene-based quantum heat engines, highlighting the role of ensemble choice and carrier-number constraints in setting efficiency limits.

### v3

@@ I Introduction
Quantum thermodynamics examines how the fundamental principles of energy transfer, heat exchange, and entropy apply to small systems where quantum effects such as coherence, entanglement, spin correlations, and discrete energy levels give rise to thermodynamical properties with no classical counterpart, and where quantum statistics provides a natural framework to describe these phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13].
@@
In this context, considerable attention has been devoted to understanding how quantum systems, such as trapped ions, superconducting qubits, quantum dots [14, 15, 16, 17, 18, 19], and low-dimensional electronic materials and their diverse behaviors can be harnessed to implement heat engines capable of producing useful work. Among quantum thermal machines, the Stirling cycle has attracted particular interest due to its potential to achieve high efficiency in regimes where thermalization and isochoric processes can be controlled with precision [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. In light of these considerations, theoretical and experimental investigations have shown that quantum heat engines based on discrete energy spectra may surpass their classical counterparts in specific regimes, particularly when the working substance exhibits a tunable energy level structure and distinctive statistical properties [33, 34].
@@
Graphene and its multilayer stacks, when subjected to perpendicular magnetic fields [35, 36, 37, 38, 39] stand out as promising platforms for quantum heat engines. Beyond their ease of fabrication and high degree of experimental control [40, 41, 42, 43, 44, 45], manifestations of their quantized spectrum are robust against temperature, since these systems exhibit quantum Hall plateaus even at room temperatures under sufficiently high magnetic fields [46] – a critical feature for practical applications. Furthermore, the large degeneracy of the Landau levels (LL) involves a macroscopic number of participating electrons, $N$, even in micron-size flakes [39] This provides a unique framework to explore quantum corrections to thermodynamic quantities stemming from the discrete Landau-level spectrum, while remaining within the thermodynamic limit. Consequently, the small-$N$ thermal fluctuations typically found in few-body systems (both classical and quantum) are suppressed, making it possible to isolate genuine quantum effects. Very importantly, when driven by a varying magnetic field $B$, these quantum systems - unlike most others - avoid the dissipation associated to Landau-Zener transitions [47, 48], as Landau levels do not cross with changes in $B$. Finally, like other two-dimensional materials, the chemical potential (or doping) can be precisely tuned via electrostatic gating, offering a robust mechanism for controlling the engine’s working fluid.
@@
Several studies have proposed graphene-based systems, including quantum dots and multilayer structures, as promising candidates for quantum engines [49, 50, 51, 52, 53, 54, 55, 56]. While these pioneering works offer valuable insights, they predominantly rely on Boltzmann statistics [57, 51, 7, 50]. By neglecting the fermionic nature of the electronic working fluid and assuming charge neutrality, the lack of quantitative accuracy of their results and the narrow focus of their setting call for a reexamination.
@@
This study examines a quantum Stirling cycle implemented in monolayer, AB-stacked bilayer, and ABC-stacked trilayer graphene under perpendicular magnetic fields as the working medium. Employing a low-energy Landau-level description and Fermi–Dirac statistics at finite doping, we correct the previous studies, expand they scope beyond the charge-neutrality point, and provide a thermodynamic analysis more closely aligned with experimentally accessible conditions [58, 59]. Within a grand-canonical framework, we evaluate key observables such as internal energy and entropy while explicitly accounting for the degeneracies specific to each multilayer configuration.
@@ II.1 Landau Level Spectrum in Multilayer Graphene
At low electronic doping, the low-energy LL spectra of mono-, AB-bilayer, and ABC-trilayer graphene in a perpendicular magnetic field (as shown in Fig. 1) can be cast into a single compact form that highlights the multilayer scalings [36, 37, 38, 35, 45].
@@
where $n$ is the LL index and $\pm$ indicate conduction and valence band LLs. The $J$-dependent cyclotron frequency $\omega_{c}^{(J)}$ encodes the microscopic parameters of each system. As we show below, Eq. (1) yields the characteristic magnetic-field scaling
@@
For the three cases addressed here, the corresponding LL energies are given by
@@
$\hbar\omega_{c}^{(1)}$ $=\sqrt{2e\hbar v_{F}^{2}B}{\,\approx 36.3\,\sqrt{B}\,\mathrm{meV},}$
@@
$\hbar\omega_{c}^{(2)}$ $=\frac{2e\hbar v_{F}^{2}B}{t_{\perp}}{\,\approx 4.77\,B\,\mathrm{meV},}$
@@
$\hbar\omega_{c}^{(3)}$ $=\frac{\left(2eB\hbar v_{F}^{2}\right)^{3/2}\sqrt{\hbar}}{t_{\perp}^{2}}{\,\approx 0.771\,B^{3/2}\,\mathrm{meV},}$
where the magnetic field $B$ is measured in tesla. Here, $v_{F}$ denotes the Fermi velocity of graphene and $t_{\perp}$ represents the interlayer hopping amplitude for each system, which characterizes the electronic coupling between adjacent graphene layers. Equations (4) to (6) establish distinct characteristic scales of the Landau-level spectra in multilayer graphene systems. These differences are ultimately manifest in the thermodynamic properties discussed next.
@@ III.1 Grand canonical formalism
In recent studies using graphene as a thermal machine, the thermodynamic behavior of graphene multilayers has been investigated through thermal states based on the canonical ensemble and Landau-level spectra [56, 51], neglecting the fermionic nature of the working fluid. In contrast, in this manuscript the thermodynamic properties of electrons in graphene multilayer systems under perpendicular magnetic fields are naturally described within the grand-canonical ensemble using Fermi–Dirac statistics, where both the temperature $T$ and the chemical potential $\mu$ serve as control parameters.
@@
Figure 2: Color map of $N_{e}$ as a function of magnetic field B[T]B\,[T] and chemical potential $\mu\$[meV] for (a) monolayer, (b) bilayer and (c) trilayered graphene system under a fixed temperature $T=50$ K.
For multilayer graphene with quantized Landau levels $\varepsilon_{n,\pm}^{(J)}$, the grand-canonical partition function can be written explicitly as
@@
$S(T,B,\mu)$ $=-k_{B}\!\!\int\limits_{-\infty}^{\infty}\!d\varepsilon\,\nu^{(J)}(\varepsilon)\,[f\ln f+$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;(1-f)\ln(1-f)].$ (13)
Both approaches are formally equivalent: the discrete-level formulation based on the grand potential and the continuum representation using the DOS combined with the Fermi–Dirac distribution, both yielding to identical thermodynamic quantities. It is worth emphasizing that the DOS representation makes explicit the role of Landau-level degeneracies, particularly the zero-energy states in all systems. We adopt the latter approach for our calculations. To guarantee numerical stability across the entire parameter space, the sums over Landau levels are truncated at a cutoff index of $n=2000$, or when the relative difference between consecutive evaluations falls below $10^{-6}$. Crucially, for the specific temperature and magnetic field regimes considered, thermal population is confined to the lowest levels, with contributions from $n>20$ being exponentially suppressed. Consequently, the effective number of participating states is limited to $n\lesssim 20$, which is fully consistent with the $\delta$-function DOS representation employed in this work [60].
@@ III.2 Energy Scales and Parametrization of the Chemical Potential
Having established the theoretical framework, we now estimate the order of magnitude of the energy separation between the ground state and the first accessible Landau excitation as the magnetic field $B$ is varied. This quantity $\hbar\omega_{c}^{(J)}$ sets the relevant energy scale for the thermodynamic analysis, since it determines the temperatures required to thermally populate excited states as the field is varied [35, 36, 45, 37].
For reference to experimental scales, we provide an estimate obtained from the magnetic length, $\ell_{B}=\sqrt{\hbar/(eB)}$ [35, 39]. By defining $N_{e}$ as the number of electrons within the magnetic length area $\pi\ell_{B}^{2}$, the sheet density can be written as
@@
Fig. 2 provides an overview of the relation between the magnetic field $B$, the chemical potential $\mu$, and the resulting particle number $N_{e}$ for mono-, bi-, and trilayer graphene at fixed temperature of $T=50$ K . In this figure, $N_{e}$ is not fixed; instead, we map the full surface $N_{e}(50,B,\mu)$ in order to illustrate how the distinct Landau–level structures of each multilayer system shape the dependence of the particle number on $\mu$ and $B$. This exploration helps visualize the scaling behaviors originating from the linear, quadratic, and cubic dispersions of the three stackings.
@@ III.3 Thermodynamic Properties: Internal Energy and Entropy of Multilayer Graphene
The next discussion focuses on the entropy to further characterize the thermodynamic response. Fig. 4 presents the temperature dependence of the entropy $S(T,B,\mu(T,B))$ for monolayer, bilayer, and trilayer graphene at different magnetic fields, with the particle number fixed at $N_{e}=15$. While all three systems exhibit a monotonic increase of entropy with temperature, notable differences arise in magnitude, curvature, and sensitivity to the magnetic field, underscoring the distinct thermodynamic responses of each multilayer configuration. In Fig. 4(a), the entropy of monolayer graphene remains negligible at low temperatures and begins to rise appreciably only above $T\simeq 25\,\mathrm{K}$, where thermal excitations populate higher Landau levels. The increase is markedly nonlinear, particularly at lower magnetic fields. This behavior reflects the relativistic Landau-level structure where stronger magnetic fields enlarge the spacing between energy levels and thereby suppress thermal activation. Consequently, higher fields substantially reduce the entropy and produce a clear separation among the curves.
@@ III.3.1 Quantum Stirling Cycle
Beyond these two classical regimes, previous studies have shown that the quantum Stirling cycle may also operate in two additional regimes that arise purely from its quantum character [63, 64, 25, 30, 65]. In the accelerator regime, both the total work and the heat released to the cold reservoir are negative, indicating that external work input is accompanied by simultaneous heating of both reservoirs. In the heater regime, all heat flows are negative and $W<0$, so that the external work is entirely dissipated as heat into both reservoirs. Importantly, these operational regimes are not exclusive to intrinsic properties of stacked graphene systems. It has been shown that twisted bilayer graphene, when employed as a Stirling engine at temperatures within 50 to 150 K, can also exhibit this diversity of thermodynamic behaviors [56].
As will be shown in later sections, the exotic thermodynamic properties of the system, which arise from the constraint of particle number conservation, can induce these additional operational regimes.
@@ IV.1 Stirling Engine Efficiency
For the bilayer graphene in Fig. 6(b), the system operates as an engine across the entire $B_{L}$ range considered, with the high magnetic field fixed at $B_{H}=3.18\,\mathrm{T}$. The thermal reservoirs are maintained at $T_{L}=45.9\,\mathrm{K}$ and $T_{H}=47.7\,\mathrm{K}$, corresponding to a Carnot efficiency of $\eta_{C}\simeq 0.04$. Under these conditions, the curves for $N_{e}=5$ and $N_{e}=15$ attain the Carnot bound at finite values of $B_{L}$, with the $N_{e}=15$ case attaining $\eta_{C}$ at lower $B_{L}$ values than $N_{e}=5$. In contrast, the $N_{e}=8$ curve remains slightly below $\eta_{C}$ over the full range, exhibiting a broad maximum just under the Carnot line at $B_{L}\simeq 0.8$[T]. This reflects a more balanced competition between the isothermal and isomagnetic contributions arising from the parametrization of the chemical potential. This intermediate behavior highlights how the bilayer spectrum, with its linear scaling in $B$, provides a broader operational window: depending on the carrier number, the cycle can either saturate the Carnot limit or remain close to it with finite work output. This behavior contrasts with results for twisted bilayer graphene studied within the canonical ensemble [56]. While such systems can reach Carnot efficiency in Otto cycles [50] and exhibit high efficiencies in quantum Stirling cycles, saturation of the Carnot bound in the Stirling cycle has not been reported in that framework. This suggests that a Fermi–Dirac description with fixed particle number enhances thermodynamic performance in graphene-based quantum heat engines, highlighting the role of ensemble choice and carrier-number constraints in setting efficiency limits.